Solve the BVP: от от OT PDE: BC: T(0.t) = (5,0) = 0 (5.1) = 0 0SX 55 Or? BC: TO and T(x,0) = 1-X
Answer #1

ANSWER:
PDE(partial differential equation):

In mathematics, a partial differential equation (PDE) is a differential equation that contains unknown multi-variable functions and their partial derivatives.

Given PDE: (partial Differential Equation) PDE: Se BC : 1 (0,4) (5,t) =0 04105 and T (2,0) = 1-X T (ait) 3T(H). X (X) af (All-tet SI = -17 (t) = Cie. xa) la cos x t (sin la The total Equation is 7(2,6) = x(Q)T(t) - la cos/x+cz sinda) sett anet [cachewe know that, Cica =k;=0 CQ Cz = ka -AE T(1, t) = K, coslae + Kasintre ey - 2.0 (cs dzett) + Ke Sinize en e + Ko sindae -ART